# envname-P envname#1

### v1

@@ 6. Related Work
Supply Chain Security: Software supply chain attacks (ohm2020backstabber, ; wang2021feasibility, ; ohm2023sok, ) such as the SolarWinds attack (martinez2021software, ) and XZ Utils backdoor (helpnetsecurityBewareBackdoor, ; openwallOsssecurityBackdoor, ) have received attention from researchers and industry. Tools like Crev (crev, ) and JFrog (jfrogClosingSupply, ) provide support for auditing dependencies and tracking associated metadata. Software bill of materials (SBOM) (muiri2019framing, ; hendrick2022state, ; xia2023empirical, ; zahan2023software, ) is a general set of techniques for providing an overview of software contents, which can be used for humans understanding; these are similar to the relational model used in our work. Tools have also been built for specific languages such as JavaScript (zimmermann2019small, ; cogo2019empirical, ; zahan2022weak, ) and Go (google-capslock, ; google-capslock-blog, ; google-capslock-talk, ). Most relevant to our work, a few techniques in this space have used formal methods and formal modeling: Koscinski and Mirakhorli (formally-modeled-cwes, ) propose a technique to formally model Common Weakness Enumerations (CWEs) using an encoding in the Alloy specification language. Software configuration management (SCM) (zeller1997unified, ) is also related; it aims to identify a configuration of dependencies that satisfies formal constraints. Our work can be thought of as a similar configuration problem to optimize for trust cost, and selecting among multiple possible configurations (different versions of different software packages) would be an interesting direction for future work.
@@
Safety and Security for Rust Code: RustSec (rustsec, ) and Lib.rs (lib-rs, ) provide databases of known vulnerabilities and libraries. Cargo Vet (cargo-vet, ), Cargo Audit (cargo-audit, ), Cargo Crev (cargo-crev, ), and Cargo Scan (cargo-scan, ; cargo-scan-paper, ) (which inspired the side effects metrics used in our tool) are tools to assist with human auditing for Rust code. Verification tools for Rust include work on formalized foundations like RustBelt (rustbelt, ), Oxide (oxide, ), BoCa (wagner2025linearity, ), Stacked/Tree Borrows (stacked-borrows, ; tree-borrows, ) and others (grannan2025place, ) and verification tools such as Verus (verus, ), Prusti (prusti, ), Kani (kani, ), and Flux (flux, ). Program analysis tools like MIRChecker (li2021mirchecker, ), MIRI (miri, ), and Rudra (bae2021rudra, ) detect memory safety and undefined behaviors in Rust code, and fuzzing tools such as SyRust (takashima2021syrust, ) and RULF (jiang2021rulf, ) use automated testing to identify bugs. Sandboxing tools (e.g., (trust, ; sandcrust, ; pkru-safe, ; xrust, ; fidelius-charm, ; galeed, ; gulmez2023friend, )) aim to protect Rust code from its unsafe subset. However, even the safe subset of Rust can be exploited, and such cases are not always detected by prior tools (hassnain2024counterexamples, ; cve-rs, ; cargo-scan-paper, ), which motivates work incorporating human factors.
@@
Trust in Software Engineering: Hamer et al. (hamer2025trusting, ) explored the impact of contributor reputation and other indicator metrics in ecosystems like Rust and GitHub. Boughton et al. (measuingTrustOSS, ) proposed a framework to decompose and quantify trust within open-source software. We used these studies to motivate the specific assumptions that we support in our minimum trust cost model. Devanbu et al. (devanbu1998techniques, ) argue for combination of trusted hardware, key management and verification tools (including, e.g., proof-carrying code (necula1997proof, )) to elevate trust, and Toone et al. (toone2003trust, ) introduce trust mediation, a framework for managing complex distributed trust relationships. Models of trust have been studied outside of computer science in a human management setting, see e.g., (mayer1995integrative, ).

### v2

@@ 6. Related Work
Software supply chain attacks (Ohm et al., 2020; Wang, 2021; Ohm and Stuke, 2023) such as the SolarWinds attack (Martínez and Durán, 2021) and XZ Utils backdoor (Zorz, 2024; Freund, 2024) have received attention from researchers and industry. Tools like Crev (development team, 2024b) and JFrog (Chin, 2023) provide support for auditing dependencies and tracking associated metadata. Software bill of materials (SBOM) (Muirí, 2019; Hendrick and Zemlin, 2022; Xia et al., 2023; Zahan et al., 2023) is a general set of techniques for providing an overview of software contents, which can be used for human understanding; these are similar to the relational model used in our work. Tools have also been built for specific languages such as JavaScript (Zimmermann et al., 2019; Cogo et al., 2019; Zahan et al., 2022) and Go (Google, 2023b, a). Especially relevant to our work, a few techniques in this space have used formal methods: Koscinski and Mirakhorli (Koscinski and Mirakhorli, 2024) propose a technique to formally model Common Weakness Enumerations (CWEs) using an encoding in the Alloy specification language. Software configuration management (SCM) (Zeller and Snelting, 1997) is also related; it aims to identify a configuration of dependencies that satisfies formal constraints. Our work can be thought of as a similar configuration problem to optimize for trust cost, and selecting among multiple possible configurations (different versions of different software packages) would be an interesting direction for future work.
Safety and security for Rust code:
@@
RustSec (Rust Secure Code Working Group, 2025) and Lib.rs (Lesiński, 2024) provide helpful databases of known vulnerabilities and libraries. Cargo Vet (Mozilla, 2024), Cargo Audit (Rust Secure Code Working Group, 2024a), Cargo Crev (development team, 2024a), and Cargo Scan (PL and PLSysSec, 2026; Zoghbi et al., 2026) (which inspired the side effects metrics used in our tool) are tools to assist with human auditing for Rust code. Formal verification for Rust includes work on foundations — like RustBelt (Jung et al., 2017), Oxide (Weiss et al., 2019), BoCa (Wagner et al., 2025), Stacked/Tree Borrows (Jung et al., 2019; Villani et al., 2025), and others (Grannan et al., 2025) — and tools such as Verus (language team, 2022), Prusti (Astrauskas et al., 2019), Kani (VanHattum et al., 2022), and Flux (Lehmann et al., 2023). Program analysis tools like MIRChecker (Li et al., 2021), Miri (Rust team, 2024), and Rudra (Bae et al., 2021) detect memory safety and undefined behaviors in Rust code, and fuzzing tools such as SyRust (Takashima et al., 2021) and RULF (Jiang et al., 2022) use automated testing to identify bugs. Sandboxing tools (e.g., (Bang et al., 2023; Lamowski et al., 2017; Kirth et al., 2022; Liu et al., 2020; Almohri and Evans, 2018; Rivera et al., 2021; Gulmez et al., 2023)) typically aim to protect Rust code from its unsafe subset. However, even the safe subset of Rust can be exploited, and this is sometimes missed by prior tools (Hassnain and Stanford, 2024; user), 2024; Zoghbi et al., 2026), which motivates work incorporating human factors.
Trust in software engineering:
@@
Hamer et al. (Hamer et al., 2025) explored the impact of contributor reputation and other indicator metrics in ecosystems like Rust and GitHub. Boughton et al. (Boughton et al., 2024) proposed a framework to decompose and quantify trust within open-source software. We used these studies to motivate the specific assumptions that we support in our minimum trust cost model. Devanbu et al. (Devanbu et al., 1998) argue for a combination of trusted hardware, key management, and verification techniques to elevate trust. Toone et al. (Toone et al., 2003) introduce trust mediation, a framework for managing complex distributed trust relationships. Outside of computer science, models of trust have been studied in other settings (e.g., (Mayer et al., 1995)).
Acknowledgements.
This research was supported in part by the US National Science Foundation under award CCF #2327338. The authors would like to thank the anonymous reviewers and Prem Devanbu for feedback leading to improvements to the paper, and Audrey Gobaco for helpful discussions and work on related projects.